# Model Lakes

### v1

@@ Abstract.
Given a set of deep learning models, it can be hard to find models appropriate to a task, understand the models, and characterize how models are different one from another. Currently, practitioners rely on manually-written documentation to understand and choose models. However, not all models have complete and reliable documentation. As the number of machine learning models increases, this issue of finding, differentiating, and understanding models is becoming more crucial. Inspired from research on data lakes, we introduce and define the concept of model lakes. We discuss fundamental research challenges in the management of large models. And we discuss what principled data management techniques can be brought to bear on the study of large model management.
@@ 1. Introduction
With the rise of AI Chatbots (ChatGPT (OpenAI, [n.d.]a), BARD (Google, [n.d.]a), etc.) and other AI-integrated apps (AI, [n.d.]a; Microsoft, [n.d.]; AI, [n.d.]b; Labs, [n.d.]b; Google, [n.d.]b; Chintagunta et al., 2021; Trummer, 2022) many organizations commit significant resources to developing Machine Learning Models, most of which are fine-tuned versions of popular models such as BERT (Devlin et al., 2019), BART (Lewis et al., 2020), GPT-3 (Brown et al., 2020), LLAMA-2 (Touvron et al., 2023), and more. Recently, OpenAI created a mechanism (OpenAI, [n.d.]b) to allow users to create custom version of ChatGPT. To support this proliferation, sharing, and reuse (finetuning and editing) of large models, many models are hosted on platforms to support the collaborative use and sharing of models such as HuggingFace (Face, 2023) and Kaggle (Kaggle, 2023).
In the data management community, we have done a lot of introspection about how machine learning and Large Language models will revolutionize data management (Halevy et al., 2023; Fernandez et al., 2023). It is true that for many important (and stubbornly hard) semantic data management problems such as data discovery and data wrangling, advanced machine learning models are out performing other approaches (Fan et al., 2023b, a; Dong et al., 2023; Li et al., 2020; Narayan et al., 2022). Exciting research challenges remain in applying these models in data management and some point out the lack of exploration of models beyond LLMs (Borisov et al., 2022). However, we do not believe this makes machine learning an existential threat to data management or a challenge to the the raison d’être of data management (Fernandez et al., 2023). Rather, we argue that the large model revolution poses some new and fascinating scientific challenges that data management researchers are, perhaps, uniquely qualified to tackle. In this vision paper, we pose the question of how can principled data management revolutionize the use and management of models, including important problems like model selection (see Example 1) and model provenance. While model management (Bernstein et al., 2000) has long been a core topic in data management, we have been slow to move beyond the management of schemas and mappings (Doan et al., 2012) to the management of large machine learning models. Today, the way in which we search for models is through 20th century keyword search over (usually manually specified) names or metadata. The way in which we understand the relationship between two models is quite primitive (for example, is one manually declared to be a fine-tuned version of another). And the study of model provenance or lineage is just emerging (Mei et al., 2022; Mu et al., 2023).
We draw on lessons learned from collaborative data sharing platforms, such as Data Hub (Bhardwaj et al., 2015), that led to the development of data lakes (Nargesian et al., 2019) as powerful repositories for the management of heterogeneous collections of data. We introduce model lakes, as a parallel to data lakes, and discuss how important innovations in data lakes, including data discovery, annotation, and version management, can (and should) be applied within model lakes and studied with the same vigor.
Example 1 ().
@@
Model Selection Problem: Consider a situation where a user wants to find a model that can summarize a legal document and simplify it in a non-technical manner. On HuggingFace (as of Jan 2024), the user finds that there are around 489,000 models uploaded and around 1500 of them have the ‘summarization’ task tag. While there are filters (trending, most likes, most downloads, model name search, and more), the user finds it hard to choose which model they should actually use. There are various concerns that the user goes through as she scrolls through different model cards (a common semi-structured form of model documentation). Is this model aware of legal jargon? Is it good at summarizing and simplifying legal documents? Is this the latest version of the model? Was this model trained on legal texts and if so which texts? What are other models that are similar to this model? Are they also trained on the same or different legal texts?
2. Models
Consider a model, , which is trained with dataset, . This model has an algorithm (function), , where . Based on the training objective provided by , can be discriminative or generative in nature. In the case of discriminative models, would learn to differentiate and classify different kinds of data instances. In the generative case, can be a sampling process that produces a distribution of new data instances. For instance, text-conditional diffusion models (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) take a text input, and produce an output that contain pixel values that form an image that matches the text description. For large language models (Touvron et al., 2023; Zheng et al., 2023; Brown et al., 2020), the output can be generated text (which may be in the format of tabular or semi-structured data), that the model creates in response to an input prompt . In general, training data can be of various types or combinations — it can be text documents, images, audio files, etc., that may be labeled or unlabeled. The full algorithm would include both ’s architecture and a pipeline for training and testing it. could be a parameterized model with any architecture, such as a Convolutional Neural Network (Cun et al., 1990) or a Transformer (Vaswani et al., 2017) or a simpler model. Note that unlike others (Fernandez et al., 2023), in our work, we consider not just LLMs, but also the worker-bee models that may be discriminative or generative, and that have been driving data science innovation since before the ChatGPT era. Moreover, recent findings from a survey on deep learning and tabular data (Borisov et al., 2022) underscore the necessity of investigating novel architectures for adoption, including diffusion or hybrid classical machine learning and deep learning models. They stress this is particularly important for handling heterogeneous collections of tabular data with better predictive performance and computational complexity. By developing and advancing a semantic model management system, the exploration of these models can be conducted in a more informed and efficient manner.
3. State-of-the-art in Model Management
Model Repositories and Model Registries. The proliferation of models has lead to the emergence of model repositories such as Civitai (Inc., 2023), HuggingFace (Face, 2023), or Kaggle (Kaggle, 2023) that openly share models. Companies, such as AWS also provide model repository products (such as the Amazon Elastic Container Registry) that can be used to share models within an organization. These platforms let users search for models by browsing through catalogs or by doing keyword search over manually created model names (and sometimes metadata). Beyond the hosting and sharing of models, a model registry provides clear naming conventions (for models and versions) and standardized, structured metadata (examples include Neptune (Labs, [n.d.]a) and AWS’ Sagemaker Model Registry (Verma et al., 2022)) and is more of a comprehensive system that enables documentation of the entire lifespan of machine learning models.
Documenting Models. Models that are deployed are usually accompanied with documentation known as model cards (Mitchell et al., 2019). This report generally contains the following categories of information.
Model Details: Information about the model such as model date, version, type, training algorithms used and other resources to learn more about the model.
Intended Use: Use cases that were imagined while developing the model and related primary users.
Factors: Summary of model performance across relevant features (like gender in people-centric models), instrumentation (like recording instrument for audio-based models), and environment (like image lighting for image-based models).
Metrics: Details of model performance measures, decision thresholds, and approaches to uncertainty and variability.
Evaluation Data: Information on dataset(s) used for quantitative analyses, the motivation behind choosing the dataset(s) and related preprocessing methods.
Training Data: Similar to information present for evaluation data section, but on data used to train the models.
Quantitative Analyses: Results of evaluating a model based on the metrics, which may be broken-down based on factors mentioned earlier. This includes unitary, i.e., model performance with respect to each factor, and intersectional results, i.e., model performance with respect to the intersection of factors.
Ethical Considerations: Information on ethical concerns such as model usage of sensitive data, intention of using the model for human life-related decisions, potential ethical risks and harms.
Caveats and Recommendations: Details of concerns or suggestions for using the model that are not covered elsewhere.
Similar to datasheets (Gebru et al., 2021), model cards are designed to guide developers in documenting models in a structured way. Model cards can be and should be augmented with information more similar to nutritional labels (Stoyanovich and Howe, 2019) that also include information about fairness and bias in the data (models). They can be further enriched with lineage and security related documentation such as adversarial attack and related defense measure policies, which are outlined in FactSheets for AI services (Arnold et al., 2019).
@@ 3. State-of-the-art in Model Management
Model Search and Discovery. In Example 1, we showcase a potential scenario for the Model Selection Problem. The current solution pipeline involves a user searching for a relevant model by naming specific models or by typing related keyworks like legal to find models that either have that word in their name or in their model card. In other words, the search relies on the model’s name and documentation. Hence, any sorting of the answer by relevance is just the relevance (prevalence) of the keywords and is not a semantic notion. Of course, this search may fail if the documentation is incomplete or inaccurate. The state-of-the-art in verifying the documentation of a model is also in its infancy.
Interpretability. Interpretability has been defined as the extraction of relevant knowledge from a model concerning relationships either contained in data or learned by a model (Murdoch et al., 2019). Interpretability is critical for model analytics to enable semantic analysis and understanding of models and their training data. Several works have surveyed techniques and research questions in this area (Singh et al., 2024; Luo and Specia, 2024; Zhao et al., 2024; Räuker et al., 2022; Doshi-Velez and Kim, 2017; Gilpin et al., 2018), which can be broadly categorized into the following areas: Local model explanations explain the sensitivity of individual output predictions to local changes in inputs using gradients (Sundararajan et al., 2017), masks (Petsiuk et al., 2018; Fong et al., 2019), local models (Ribeiro et al., 2016), or Shapley values (Lundberg and Lee, 2017; Chen et al., 2023). Global explanations explain the mechanisms of the overall model at the level of attention patterns (Clark et al., 2019), representations (Belinkov, 2022; Zou et al., 2023), circuits (Elhage et al., 2021; Wang et al., 2022), or neurons (Bau et al., 2018, 2017; Geva et al., 2023). Models can be designed to be inherently interpretable (Rudin et al., 2022). Lastly, datasets can be explained using natural language explanations (Narayan et al., 2022), data visualizations (Dibia, 2023), or by training inherently interpretable models (Ustun and Rudin, 2016; Singh et al., 2023) of the data set.
Provenance. A new and important line of inquiry considers questions related to provenance (Mei et al., 2022; Mu et al., 2023). These issues pose similar questions to those studied in the database community (Buneman and Tan, 2018): from where did a generated fact derive or why was a predication made? A subset of these questions are also posed in model interpretability as part of local and global explanations. As in data provenance, the main issue is in whittling down all the provenance associated with some process (such as a model) into simple, but useful explanations (Buneman and Tan, 2018). More than two decades of research in the data management community have produced elegant and simple, yet powerful, models for data provenance when the process is a query (Buneman et al., 2008, 2016) or workflow (Moreau et al., 2013; Moreau and Groth, 2013). In large model provenance, the goal cannot be to understand and track all inputs (data, hyperparameters, code, training regimes, …) that were used to create a specific model output. Rather research is needed on finding meaningful sets of concepts that can be tracked efficiently and which provide important insights into model behavior.
There has been some initial research on solving the model provenance. Mu et al (Mu et al., 2023), for instance, uses a data-driven and model-driven representation learning method to encode ‘Model DNA’, which the authors then use as part of a framework for model provenance identification. Their goal is identify whether a model is a pre-trained model of another model. The former model is known as the source model while the latter model is known as the target model. This technique assumes a relaxed condition where the source and target model have the same architecture or structure and the target model’s training is based entirely on the source model. Furthermore, it assumes that we can access training data of the source model. However, there are harder conditions that can exist where the target model is trained based on a subset of the parameters of the source model or when the architecture of the target model is different from that of the source model. We define the granularity of such model conditions in Section 5 and explore various ways a model can be different from one another.
Model Versions. Models are valuable assets that are reused in numerous ways to create new versions. Based on the training type, the function or algorithm, can vary. For instance, if a model is pre-trained, would include a pipeline were it would encourage to learn general features from dataset . However, if is fine-tuned, would start off with weights from a previously trained model and learn features from to understand a specific task better. While pre-training and fine-tuning are some of the most common training objectives for creating new versions of a model, there are newer, more surgical-like training objectives such as model editing (Sinitsin et al., 2020; Meng et al., 2022) where focuses on updating certain facts (for instance, updating the name of the current President of a country) learned by a model in a more localized fashion. With various options for training methods, architecture, and dataset combinations, enormous numbers of models have been created with varying degrees of similarities and differences amongst each other.
Among the versioning methods, we examine fine-tuning and editing in more detail. Fine-tuning is the process of further training the parameters of an existing model to improve its performance, usually on specific task(s) or domain(s). For instance, T5 (Raffel et al., 2020) is a pre-trained model that has been trained on a collection of large collection of text (about 750 GB) to perform well on a diverse set of tasks. This has been further fine-tuned with the MIMIC-III (Johnson et al., 2016) and IV (Johnson et al., 2023) dataset to form Clinical-T5 (Lu et al., 2022) to perform better for medical domain-related tasks. Another model instance is FLAN-T5 (Chung et al., 2022) that has been fine-tuned on data with 1.8K tasks phrased as instruction so that it can handle such prompts better. Apart from datasets and training goals, there are also various optimizations schemes that could differentiate various fine-tuned models. For instance, Hu et al. (2022) experimented with a GPT-2 model fine-tuned with Adam (Kingma and Ba, 2015) and another GPT-2 model instance fine-tuned with their Low Rank Adaption (LoRA) (Hu et al., 2022). They both have the same task goals and are fine-tuned from the same source model, but have a difference in performance due to the fine-tuning optimization conducted. Consequently, with different datasets, training goals and algorithms, various versions of a pre-trained model can be created. Because they have the same source model, the fine-tuned models can also be considered to be related to each other.
Direct model Editing. Large language models (LLMs) contain a wealth of knowledge encoded in their parameters. As a result, they can respond to natural-language queries such as those that relate to factual statements. However, LLMs can sometimes hallucinate or point to outdated information. Hence, there have been several studies that look into localizing and understanding the internal mechanisms within LLMs. Zhu et al. (2020), for instance, proposed the task of modifying factual knowledge in transformer models such as BERT (Devlin et al., 2019) by performing a constrained fine-tuning on the original model with specific facts. De Cao et al. (2021), on the other hand, introduced Knowledge Editor, that uses hyper-networks (Ha et al., 2017) to tune the model’s parameters for editing specific knowledge. In terms of finding these facts, Dai et al. (2022) proposed the idea of knowledge neurons, whose activation highly correlated with their related knowledge expression. MEND (Mitchell et al., 2022) also edits local knowledge, but it does so by training a collection of Multi-Layer Perceptrons (MLP) to modify model gradients that would result in the production of local model edits. In contrast, ROME (Meng et al., 2022) uses rank one modification of the MLP weights of a single layer to directly edit appropriate weights that would change the fact associated to an information present the transformer model. MEMIT (Meng et al., 2023) builds upon ROME to be able to edit multiple knowledge expressions in the model. Direct model editing can also be applied to multimodal models: the TIME (Orgad et al., 2023) and UCE (Gandikota et al., 2024) methods edit text-to-image diffusion models. All these methods showcase various ways to edit a model, and hence, can be useful to understand how a current model is changed from its base model when a fact is changed.
4. Model Lakes: Vision
Given this state-of-the-art, we now outline our vision for model lakes: powerful repositories for the management of heterogeneous models. We enumerate challenges that require new scientific solutions and provide evidence that these challenges can be met.
Content-based Model Search. In analogy with dataset search, Google provides a comparable service to model registries, but for datasets (Benjelloun et al., 2020; Alrashed et al., 2021). They require dataset owners to follow their naming conventions and use metadata from Schema.org. Note this type of search only works well when there are clear incentives for artifact owners to provide complete and compliant metadata. Google provides this by returning datasets that comply with its best practices, though even Google struggles with impostors who game the system (Alrashed et al., 2021) by giving invalid semantic annotations. Current model repositories, such as Hugging Face, have found that they are subject to model designers who give sparse or ambiguous metadata within their model names or model cards (Liang et al., 2024). There is a danger that people could intentionally misinform with malicious intent (Security, [n.d.]b).
@@ 4. Model Lakes: Vision
Within open data and enterprise data, researchers have learned that they cannot rely on metadata for datasets to be accurate, complete, or consistent within a data lake (Nargesian et al., 2019). As a result, there is a great deal of work on content-based search in data lakes including join search (Zhu et al., 2016, 2019; Khatiwada et al., 2022), union search (Nargesian et al., 2018; Khatiwada et al., 2023; Fan et al., 2023b), and related dataset search (Fernandez et al., 2018; Chapman et al., 2020). As mentioned in the introduction, machine learning models have revolutionized content-based dataset search. Important and impactful work has shown how we can use machine learning to create meaningful semantic representations of tuples, columns, and full tables to enhance dataset search and other semantic tasks like data integration and alignment.
But what about content-based model search? There is a bit of “physician, heal thyself” here. To the best of our knowledge, this is an area in its infancy. Lu et al. (2023) are the first to introduce the task of content-based model search, ”which aims to find the most relevant deep image generative models that satisfy a user’s input query”. Their solution builds a representation of the output distribution of a generative image model and uses this to find models that are most relevant to an input image or to an input model (or more precisely to its distribution). The trend of using meta-learning (and creating a hyper-network) for this task continues in other methods, such as AutoMRM, which is a model retrieval method based on multi-modal query and meta-learning (Li et al., 2023). A current limitation of these methods are that they are only able to perform retrieval for models that generate images. HuggingGPT (Shen et al., 2024) overcomes this limitation by using an LLM (in their case, ChatGPT) as a controller to decide which models to use based on a user’s prompt. This, however, is a metadata-based search and not a content based one, since the LLM would parse the user’s prompt into solvable tasks and then utilize another input of model descriptions (metadata) to decide which models would be relevant for each task. This approach allows queries and solutions in any modality or in any domain. However, it is limited by the capabilities of the LLM employed and the quality and accuracy of the model documentation. Furthermore, unlike content-based search, metadata-based ones would primarily fail when object-centric tags do not describe the model fully. As part of our vision, it is crucial that the content-based model search methods cover all types of models in the model lakes, including large language models. Additionally, ensuring their commercial usability, driven mainly by speed and accuracy, is essential. A hybrid approach combining metadata and content-based search could be effective, especially when exhaustive and verified metadata is accessible.

### v2

@@ Abstract.
Given a set of deep learning models, it can be hard to find models appropriate to a task, understand the models, and characterize how models are different one from another. Currently, practitioners rely on manually-written documentation to understand and choose models. However, not all models have complete and reliable documentation. As the number of models increases, the challenges of finding, differentiating, and understanding models become increasingly crucial. Inspired from research on data lakes, we introduce the concept of model lakes. We formalize key model lake tasks, including model attribution, versioning, search, and benchmarking, and discuss fundamental research challenges in the management of large models. We also explore what data management techniques can be brought to bear on the study of large model management.
@@ 1. Introduction
With the dramatic rise in AI capabilities across a variety of domains (Deng et al., 2023; Schick et al., 2024; OpenAI, [n.d.]; Google, [n.d.]a; AI, [n.d.]a; Microsoft, [n.d.]; AI, [n.d.]b; Labs, [n.d.]b; Google, [n.d.]b; Trummer, 2022; Chintagunta et al., 2021), many organizations have begun to commit significant resources to developing Machine Learning Models. Many of these are fine-tuned versions of popular foundation models, such as Llama-3 (Wang et al., 2024b), Mistral (Jiang et al., 2023), DeepSeek-R1 (DeepSeek-AI et al., 2025), Stable Diffusion (Rombach et al., 2022), BART (Lewis et al., 2020), and BERT (Devlin et al., 2019). Proprietary closed-source models such as GPT-4 (OpenAI, 2023), Gemini (Anil et al., 2023) and Claude-3 (Anthropic, 2024) also support creation of fine-tuned models. To support this proliferation, sharing, and reuse of large models, many models are hosted on platforms to support the collaborative use and sharing of models such as Hugging Face (Face, 2023) and Kaggle (Kaggle, 2023).
As the number of pre-trained models grows, comparing them and selecting the right one for specific tasks becomes increasingly difficult (see Example 1.1). Documentation, particularly model cards (Mitchell et al., 2019), aims to provide essential insights, but Liang et al. (2024) have revealed a concerning lack of transparency and completeness in these resources. This makes it hard for users to make informed decisions, especially when navigating model sharing platforms. Efforts such as the Data Provenance Initiative (Longpre et al., 2023), MLCommons (Reddi et al., 2019; Mattson et al., 2019), and Responsible Foundation Model Development Cheatsheet (Longpre et al., 2024) have been introduced to enhance the documentation of model creation and capabilities, with a particular focus on detailing their training data. However, not all model creators adhere to these guidelines, meaning that many existing and future models may still lack crucial information needed by users. To address this gap, we propose a systematic breakdown and formalization of tasks for “model lakes" — a system containing numerous heterogeneous pre-trained models and related data in their natural formats (one example is Hugging Face (Face, 2023)).
We introduce model lakes, as a parallel to data lakes, and discuss how important innovations in data lakes (Nargesian et al., 2019), including data discovery, annotation, and version management, can (and should) be applied within model lakes and studied with the same vigor. We look at how model lakes are currently managed and define and discuss tasks that can be used to better inform users about the models and their relationships. Others have discussed how one type of model (LLMs) may disrupt data management (Fernandez et al., 2023). We focus on how data management can transform the management and use of AI models.
Example 1.1.
@@
Model Search Problem: Consider a situation where a user wants to find a model that can summarize a legal document and simplify it in a non-technical manner. On Hugging Face (as of Sept 2024), the user finds that there are around 1M+ models uploaded and 1950 of them have the ‘summarization’ task tag. While there are filters (trending, most likes, most downloads, model name search, and more), the user finds it hard to choose which model to use. There are various concerns that the user goes through as she scrolls through different model cards (a common semi-structured form of model documentation). Is this model aware of legal jargon? Is it good at summarizing and simplifying legal documents? Is this the latest version of the model? Was this model trained on legal texts and if so which texts? What are other models that are similar to this model? Are they also trained on the same or different legal texts?
2. Three Viewpoints of a Model
An AI model, , can be analyzed from three viewpoints: according to its history, intrinsic composition, or extrinsic behavior. These viewpoints highlight different aspects of the model’s characteristics, aspect that we will show are useful in analyzing model lake problems and solutions.
The history of the model is defined by its training data () and training algorithm (), which may include processes like fine-tuning, model editing, or other adaptation techniques. The degree to which history is documented within a model lake can vary greatly (Liang et al., 2024).
The intrinsic viewpoint concerns the model’s internal structure. This includes the model architecture (), and the specific trained parameters (). The architecture refers to the structural design of the model, such as a combination of multi-layer perceptrons (MLPs) and attention layers in transformer models. Formally, the architecture defines a function family which is instantiated with the specific learned parameters to create the function
In contrast, the extrinsic viewpoint focuses on the model’s observable behavior and performance in user-defined tasks. The model parameters, , are not visible as part of the extrinsics. However, the function and the model behavior, , is extrinsic. For example, in the case of an unconditional generative model, the extrinsics correspond to the observable probability distribution defined by the model, . For a classifier, the extrinsics are defined by the behavior of the modeled predictions . Extrinsics can be observed in terms of a neural network’s action on inputs and outputs and , without requiring any knowledge of its training data or its internal structure.
The distinction among intrinsics, extrinsics, and history is useful because, while every model has all these characteristics, there are cases where certain aspects may be unavailable. For example, in a Model Lake, the intrinsic details of some models might be inaccessible, and some analysis methods may rely solely on extrinsic observations or historical records to understand a model’s behavior. We use this distinction to analyze possible solutions to a variety of model lake tasks. Our envisioned model lake is depicted in Figure 1.
3. Formalizing Model Lake Tasks
Model lake tasks are specifically concerned with gathering and presenting insights about the models that are stored within the lake. Topics related to the infrastructure behind model training or storage systems fall outside this scope.
Model Attribution. In data lakes, data provenance is the “description of the origins of a piece of data and the process by which it arrived in a database" (Buneman et al., 2001). Model provenance (often called attribution) considers questions like “Why was image X generated when the model was given input Y?" If the history is recorded, the history can be consulted to directly ask the training data attribution question formally: which training data items are most influential on the decision; in other words, which , if they were not present in the training data, would cause the decision to change the most? When history is not available, attribution can be studied using intrinsic and/or extrinsic clues to provide insight into the attribution of model decision behavior. For example, we can perform sensitivity analysis on the observable extrinsics of a model by asking: which aspects of the inputs to or are most important in a model’s prediction of a particular output? And if we have access to model intrinsics, we can study feature or representation analysis, which asks, which internal representations or internal “concepts” within the model are most important for a decision?
Model Versioning. In the model versioning task, we want to understand whether (and possibly how) a model has been created from other models (similar to studying how a version of a data set or table may have been derived from another (Shraga and Miller, 2023)). One possible definition of model versioning that uses an intrinsic viewpoint (Horwitz et al., 2024) is: Given a model, and a set of models, , construct a directed Model Graph, , where a directed edge between models indicates that one model is a version of the other. The edges can describe the transformation. This can include training techniques, optimization techniques, as well as the data used to further train (fine-tune) models. An important problem in versioning is: given a model’s training parameters and a candidate model’s parameters, , is a source of ? Being a source model would mean that or a version of was used for training .
Model Search. Model search is the task of finding a related or desired model. Again we can consider this task from different viewpoints. An extrinsic view considers the behavior of the model. Given a task function, where a task takes an input, and produces an output , we want to find the best performing model, , for the given task among a set of candidate models, . Each candidate model is characterized by their observable behavior, i.e., . The optimal model is selected based on a scoring function of how close the model’s behavior is to a query model.
Even considering only an extrinsic view, there are many formulations of model search. If the goal is to find models that perform similarly on specific data (for example, a specific image), then a possible definition is (Lu et al., 2023): given a point, , and a set of candidate models, , rank models based on their similarity of their respective observable behavior w.r.t. to . This could be extended to other types of models where for a given model, and a set of candidate models, , we want to find related models based on a ranking function that considers the similarity of models’ output distribution and semantic concepts/patterns present within them.
An intrinsic view of model search would find models with similar model architectures and training parameters. Or as done in data lakes where intrinisic search is the norm, we could create embeddings representing the important features of the model and design a fast nearest neighbor search over these embeddings. Considering model history, we could search for models that have been trained on the same dataset. This is straight-forward when history is recorded, but when it is not fully explicit, we may leverage extrinsic or intrinsic clues in the search.
Benchmarking. For single model tasks, a benchmark, (for example, a set of images), is used to measure the performance of a model, (or set of models) based on a scoring function, . For model lake tasks, we will need new (shared) model lake benchmarks (large sets of models ) that can be used to measure the quality of a lake task solution. This means that within a benchmark lake, we will need verified ground truth.
4. State-of-the-art in Model Mgmt
Model Repositories, Registries, and Lakes. A model repository is a storage system for machine learning models. A model registry goes beyond basic storage by offering version control (the representation of versions, not today the discovery of version relationships mentioned in Section 4). These registries typically enforce clear naming conventions (for models and versions) and organize structured, standardized metadata (Labs, [n.d.]a; Verma et al., 2022). Recent advancements have added functionalities like explainable network intrusion detection (Spadari et al., 2024) for security and integrations of different tools such as model monitoring and experiment tracking into a single framework (Vishwambari and Agrawal, 2023). Model registries are typically private, facilitating collaboration within organizations, while open platforms such as Civitai (Inc., 2023), Hugging Face (Face, 2023), and Kaggle (Kaggle, 2023) enable public sharing of models. These platforms, which we can consider as model lakes, assist users in model exploration by providing curated catalogs and keyword search, often leveraging both manually created model names and metadata for more efficient discovery. Unlike lifecycle creation and training platforms, model lakes focus on managing a set of AI models, including their interrelationships. This is distinct from model management in databases, which deals with schemas and their mappings (Bernstein and Melnik, 2007). Current model lakes capabilities lack effective mechanisms for representing and navigating the model space semantically, particularly when model documentations or names are incomplete or unknown to users, leaving many models effectively undiscoverable (Liang et al., 2024).
Documenting Models. Models that are deployed are usually accompanied with documentation known as model cards (Mitchell et al., 2019). Model cards contain (among other categories) information on model details, intended use, metrics, training data, and quantitative analyses. Similar to datasheets (Gebru et al., 2021), model cards are designed to guide developers in documenting models in a structured way. Model cards can be and should be augmented with information more similar to nutritional labels (Stoyanovich and Howe, 2019) that also include information about fairness and bias in the data (models). They can be further enriched with lineage and security related documentation (Golpayegani et al., 2024) such as adversarial attack and related defense measure policies, which are outlined in FactSheets for AI services (Arnold et al., 2019).In addition to research on enhancing the completeness of model documentation, there remains a critical gap in the verification of model cards. There is a danger that people could intentionally misinform model users with malicious intent (Security, [n.d.]b). The state-of-the-art in verifying the documentation of a model is notably in its infancy (Longpre et al., 2024). Mithril AICert (Security, [n.d.]a) is developing a certification initiative that verifies whether a model was trained using the specified algorithms and data sets. However, this initiative has some limitations. First, as it is still in development, it is not yet available for production use. Second, the certification process depends on the voluntary participation of the model creators. Furthermore, the AICert website (Security, [n.d.]c) highlights several critical limitations, including the inability to audit training code or data for risks such as backdoors or data poisoning.
@@ 4. State-of-the-art in Model Mgmt
Model Search and Discovery. In Example 1.1, we present a potential scenario for the Model Search Problem. The current solution pipeline involves a user searching for a relevant model by naming specific models or by typing related keywords like legal to find models that either have that word in their name or in their model card. In other words, the search relies on the model’s name and documentation. Hence, any sorting of the answer by relevance is just the relevance (prevalence) of the keywords and is not a semantic notion based on the model itself. Of course, this search may fail if the documentation is incomplete or inaccurate.
@@
Within open data and enterprise data, researchers have learned that they cannot rely on metadata for datasets to be accurate, complete, or consistent within a data lake (Nargesian et al., 2019). As a result, there is a great deal of work on semantic or content-based search in data lakes including join search (Zhu et al., 2016, 2019; Khatiwada et al., 2022; Dong et al., 2023), union search (Nargesian et al., 2018; Khatiwada et al., 2023; Fan et al., 2023b; Hu et al., 2023), and related dataset search (Fernandez et al., 2018; Bogatu et al., 2020). Machine learning models have revolutionized content-based dataset search. Important and impactful work has shown how we can use machine learning to create meaningful semantic representations of tuples, columns, and full tables to enhance dataset search and other semantic tasks like data integration and alignment.
But what about content-based model search? To the best of our knowledge, this is an area within ML that is in its infancy. Recent techniques for image models leverage meta learning (Lu et al., 2023; Li et al., 2023). HuggingGPT (Shen et al., 2024) uses an LLM (in their case, ChatGPT) as a controller to decide which models to use based on a user’s prompt. This metadata-based search differs from content-based approaches, as the LLM parses the user’s prompt into tasks and uses model descriptions to select relevant models. While it allows queries across any modality or domain, it is limited by the LLM’s capabilities and the quality of model documentation. Additionally, it may fail when object-centric tags do not fully capture the model. Our vision emphasizes that content-based model search must cover all models in model lakes, including large language models, while ensuring usability through speed and accuracy.